# Form and meaning co-determine the realization of tone in Taiwan Mandarin spontaneous speech: the case of T2-T3 and T3-T3 tone sandhi

### v1

@@ 2.1 Data
Data analysed in the present paper is taken from the Taiwanese Mandarin Spontaneous Speech Corpus (Fon, 2004), which consists of 30 hours of spontaneous conversations from speakers aged between 20 and 60 (31 females and 24 males). From this corpus we extracted 3937 word token with the T2-T3 tone pattern, representing 299 word types, and 4791 word tokens with the T3-T3 tone pattern, representing 269 word types. F0 measurements were obtained using Praat (Boersma and Weenink, 2020).
To ensure that for statistical analysis, sufficient numbers of tokens are available for each word type, words with fewer than 12 tokens were excluded from our dataset. For high-frequency words with more than 200 tokens, we randomly sampled 200 tokens, in order to avoid model predictions from being biased towards high-frequency words (Gahl and Baayen, 2022). As a consequence, every word type has a maximum of 200 tokens and a minimum of 12 tokens. Furthermore, in order to facilitate the statistical analysis of effects of gender, words contributed by only female speakers or only male speakers were excluded. We also made sure that tokens of one word type are contributed by two or more speakers, so as to avoid model prediction being biased by one specific speaker’s way of speaking. Tokens with potential f0 tracking errors were removed on the basis of visual inspection. The resulting dataset contains 1462 tokens representing 20 word types with tone pattern T2-T3 and 878 tokens representing 20 word types with tone pattern T3-T3.
@@ 2.2 Predictors
In our statistical analyses, the response variable is log-transformed pitch. The log-transformation was motivated by a Box-Cox analysis (Box and Cox, 1964). We modelled pitch as a function of normalized time. Table 1 provides an overview of variables that we considered in the statistical analysis.
Table 1: Predictors in the statistical analysis
Abbreviation in R Variable Definition
normalized_t normalized time For each token, time was rescaled between 0 and 1. Speech rate was included as covariate to bring effects of durational differences under statistical control.
gender gender Speakers were identified as male or female.
@@
tone_pattern tonal pattern Rising-dipping (23) or dipping-dipping (33).
spoken_freq corpus frequency Number of occurrences of a word type in the Taiwan corpus.
written_freq word frequency Number of occurrences of a word type in the Sinica Corpus of Taiwan Mandarin (Ma et al., 2001).
tonal_context tonal context A factor with 36 levels representing the combinations of the tone of the syllable immediately preceding a token and the tone of the syllable immediately following that token. Immediately preceding or following pauses were coded with PAUSE instead of a tone specification.
speech_rate speaking rate Local speech rate, defined for a token as the number of syllables per second calculated over the time period of four characters to the left and four characters to the right of that token.
@@
norm_utt_pos word position in utterance The normalized position of a token in its utterance, represented on a scale from 0 to 1. Words in single-word utterances were coded with 1.
@@
bg_prob_prev bigram probability of the preceding word The probability of the occurrence of the target word given the preceding word (following (Gahl, 2008)).
bg_prob_fol bigram probability of the following word The probability of the occurrence of the following word given the target word.
speaker speaker A factor with anonymized speaker identifiers as levels.
@@
word word A factor with orthographic word, as available in the corpus, as levels. In this study, all words under investigation are two-character words (e.g., 媒體 (mei2ti3, media).
@@
sense sense A factor with word sense instead of orthographic word, as levels.
genderXtone genderXtone A factor for the interaction of gender by tone_pattern.
Chuang et al. (2024) reported that word type is a strong predictor of tonal realization, outperforming a range of segmental properties of words’ forms such as vowel height, onset, and rhyme structure. In the present study, we therefore use word as critical predictor, avoiding issues of collinearity that come with adding in segmental properties as predictors.
The study by Chuang et al. (2024) also argues that the effect of word type is semantic in nature. In support of this hypothesis, they show that prediction of pitch contours improves when (orthographic) word is replaced by word sense. Therefore, we made use of the same word sense identification system as used in their study, described in Hsieh et al. (2024), which uses BERT in combination with the Chinese WordNet (Huang et al., 2010).  [^1] The authors thank Yu-Hsiang Tseng for identifying word sense for the dataset in the current paper.
In our dataset, there are 4 words which have no word sense in the Chinese WordNet. For the remaining 36 words in our dataset, a total of 71 senses was identified. Most of the words have 1 to 3 senses. The distribution of senses is skewed towards the right, with about one-third of the senses having 6 tokens or fewer. To make sure that all senses have sufficient tokens for statistical evaluation, we first removed senses with fewer than 8 tokens. In addition, to prevent model predictions from being biased in favor senses with a large number of tokens, we sampled 40 tokens for each of the more frequently used senses. The resulting second dataset, prepared specifically for assessing the role of sense as opposed to orthographic word, contains in total 1144 tokens of 46 sense types (corresponding to 35 orthographic word types).
@@ 2.3 Statistical analysis
One of the major challenges in analysing dynamic speech data, such as pitch contours, formant trajectories and tongue movements, is non-linear change over time (Wieling et al., 2016; Wieling, 2018; Chuang et al., 2021). The Generalized Additive Mixed Model (GAMM, Wood, 2017) has emerged as a flexible and powerful statistical method for analysing non-linear time-series data. It has been successfully applied to the analysis of f0 contours in various languages, including Mandarin (Chuang et al., 2024), Brazilian Portuguese (da Silva Miranda et al., 2020), Papuan Malay (Kaland et al., 2023), and French (Deng et al., 2023).
In this study, we make use of the mgcv package (Wood, 2017) for R (Team, 2020) , using the bam function to model pitch as a function of time and the predictors listed above. As the distribution of pitch values has heavy tails, following Chuang et al. (2024), we use a scaled-t model for the errors. According to this model, $(y-\mu)/\sigma\sim t_{\nu}$, with $\sigma$ and $\nu$ parameters that are estimated alongside the smoothing parameters of the model. In the model specification reported in the next section, the directive family=’scat’ requests the scaled-t model.
We made use of thin plate regression spline smooths to model the effect of time and its interactions with gender, tone pattern. Thin plate regression splines were also used for taking into account the bigram probabilities of preceding and following words. We used tensor product smooths for modeling interactions of time and continuous predictors such as speech rate, utterance position, and bigram probabilities. We used the ti() directive to single out the effect of the interaction and tease it apart from its corresponding main effects. Factor smooths for time were incorporated for speaker, and separately, for word. These factor smooths (specified as s(time, pred, bs="fs", m=1)) represent nonlinear random effects of subject and word, and include shrunk estimates of by-subject and by-word intercepts.
@@ 3 Results
The model that we proposed for the present data was obtained by fitting a series of increasingly complex GAMM to the dataset. The best model resulting from this exploratory investigation of the Mandarin pitch contours was specified follows:
pitch~genderXtone+

### v2

@@ 2.1 Data
Data analyzed in the present paper is taken from the Taiwanese Mandarin Spontaneous Speech Corpus [16], which consists of 30 hours of spontaneous conversations from speakers aged between 20 and 60 (31 females and 24 males). The audio recordings were orthographically transcribed using traditional Chinese characters. Word and character boundaries in the orthographic representation were identified using the word segmentation system developed by Academia Sinica [33]. The transcriptions written in Chinese characters were then romanized. The audio and transcripts were forced-aligned at both the word and character level with EasyAlign [19]. The resulting alignments were manually checked and corrected by native speakers when necessary. In the current study, we followed the segmentation provided in the corpus. The Hanyu Pinyin transcriptions and canonical tones were obtained from the pinyin_jyutping_sentence package for Python. For Chinese words with multiple possible pronunciations, the tonal specifications were verified and corrected when necessary (e.g., 东西 can be pronounced either as dong1xi0 `thing' or as dong1xi1 `east and west').
F0 values were estimated using Praat’s To Pitch (cc) function with the time step set at 0.001 seconds and a Gaussian window for optimal accuracy [1]. Pitch floor was set at 75 Hz and pitch ceiling at 400 Hz for female speakers, and pitch floor at 50 Hz and pitch ceiling at 300 Hz for male speakers. No F0 values were interpolated for the voiceless sections of the tokens.
We extracted 3937 word tokens with the T2-T3 tone pattern, representing 299 word types, and 4791 word tokens with the T3-T3 tone pattern, representing 269 word types. To ensure that for statistical analysis, sufficient numbers of tokens are available for each word type, words with fewer than 12 tokens were excluded from our dataset. Following [9], for high-frequency words with more than 200 tokens, we randomly sampled 200 tokens, in order to avoid model predictions from being biased towards high-frequency words. This led to the exclusion of 1400 tokens of 7 high-frequent word types. As a consequence, every word type has a maximum of 200 tokens and a minimum of 12 tokens. Furthermore, in order to facilitate the statistical analysis of effects of gender, words contributed by only female speakers or only male speakers were excluded. We also made sure that tokens of one word type are contributed by two or more speakers, so as to avoid model prediction being biased by one specific speaker’s way of speaking. Tokens with potential F0 tracking errors were removed on the basis of visual inspection. Table 1 presents an overview of tokens and word types after data sampling. The first resulting dataset contains 1462 tokens representing 20 word types with tone pattern T2-T3 and 878 tokens representing 20 word types with tone pattern T3-T3 (see Table 1 for the dataset of words).
The study by [9] argues that semantics co-determines the tonal realizations. In support of this hypothesis, they show that prediction of pitch contours improves when (orthographic) word is replaced by word sense. Therefore, we made use of the same word sense identification system as used in their study, described in [22], which uses BERT in combination with the Chinese WordNet [23].
In our dataset, there are four words which have no word sense in the Chinese WordNet. For the remaining 36 words in our dataset, a total of 71 senses was identified. Most of the words have 1 to 3 senses. The distribution of senses is skewed towards the right, with about one-third of the senses having 6 tokens or fewer. To make sure that all senses have sufficient tokens for statistical evaluation, we first removed senses with fewer than 8 tokens, which resulted in the exclusion of one word type. In addition, to prevent model predictions from being biased in favor senses with a large number of tokens, we sampled 40 tokens for each of the more frequently used senses. The second resulting dataset, prepared specifically for assessing the role of sense as opposed to orthographic word, contains in total 1144 tokens of 46 sense types, corresponding to 35 orthographic word types (see the sense dataset in Table 1).
Table 1: Summary of tokens and word types in each tone pattern in the word and sense analyses.
Word dataset Sense dataset
Tone pattern Tokens Word types Tokens Word types Sense types
T2-T3 1462 20 567 16 20
T3-T3 878 20 577 19 26
Total 2340 40 1144 35 46
@@ 2.2 Predictors
In our statistical analyses, the response variable is log-transformed F0. The log-transformation was motivated by a Box-Cox analysis [2]. We modelled log-transformed pitch as a function of normalized time.
Time is an important factor co-determining the realization of pitch contours, which is why token duration and its interaction with normalized time are included in the GAMM models. In the current study, however, we are interested in the shape of pitch contours, for which working in normalized time is essential. For example, consider a sine wave with some fixed amplitude, but one with a period of $2\pi$ and the other with a period of $4\pi$. Over the interval $0,2\pi]$, the curves appear dissimilar, but when normalized to the [0,1], it is clear that their shapes are identical. Since a word's duration can be predicted using a separate model [[, see, e.g.,]]gahl_time_2024, shape and duration models jointly can account for observed pitch contours in actual real time.
Time normalization was carried out on entire words, including onset consonants11endnote: 1For detailed discussion of the way in which GAMs deal in a principled way with voiceless initial or medial obstruents, see [9]. and final nasals if present. Table 2 provides an overview of variables that we considered in the statistical analysis.
Table 2: Predictors used in the statistical analysis
Variable (Abbreviation in R) Definition
Log-transformed pitch (logF0) Natural logarithm (loge) of the fundamental frequency (F0).
Normalized time (normalized_t) For each token, time was normalized to a range between 0 and 1, allowing tokens with varying durations to be modeled on a common scale.
Gender (gender) A factor of two levels: male and female.
@@
Tone pattern (tone_pattern) The tonal pattern of the token, represented as a factor of two levels: 23 (rising-dipping) and 33 (dipping-dipping).
GenderXtone (genderXtone) A factor for the interaction between gender and tone_pattern, with four levels: female.23, male.23, female.33, and male.33. This factor is used to account for gender-related differences in tone realization.
Duration (duration) The duration of the token in seconds. Duration was included as a covariate to control for the effect of time in tonal realization.
@@
Word position in utterance (norm_utt_pos) The normalized position of a token within its utterance, ranging from 0 to 1. Smaller values indicate earlier positions, and larger values indicate later positions in the utterance. Words in single-word utterances were coded as 1.
@@
Bigram probability of the preceding word (bg_prob_prev) Probability of the target word given the preceding word (following [18]), measuring how predictable the word is in context. Higher values indicate greater predictability.
Bigram probability of the following word (bg_prob_fol) Probability of the following word given the target word.
Tonal context (tonal_context) The interaction of the preceding and following tones of a token. Each tone (or a pause) can take six values: 0, 1, 2, 3, 4, or PAUSE, resulting in 36 possible tonal contexts.
Speaker (speaker) A factor with 55 anonymized speaker identifiers.
@@
Word (word) A factor with 40 orthographic word forms. All words in this study are two-character words (e.g., 媒体 (mei2ti3, `media').
@@
Sense (sense_type) A factor with 46 levels representing word senses rather than orthographic forms.
@@ 2.3 Statistical Analysis
One of the major challenges in analyzing dynamic speech data, such as pitch contours, formant trajectories and tongue movements, is non-linear change over time [50, 51, 10]. The Generalized Additive Mixed Model [[, GAMM,]]wood_generalized_2017 has emerged as a flexible and powerful statistical method for analysing non-linear time-series data. It has been successfully applied to the analysis of F0 contours in various languages, including Mandarin [9], Brazilian Portuguese [34], Papuan Malay [25], and French [13].
In the present study, we make use of the mgcv package [52] for R [44] , using the bam function to model pitch as a function of time and the predictors listed above. To account for autocorrelation in the residuals, we include a first-order autoregressive correlation (AR(1)) in the model. This assumes that the residual at time $t$ is correlated with the residual at time t1t-1, allowing the model to appropriately handle temporal dependence in the pitch contours. An AR(1) with $\rho=0.89$ removes most of the autocorrelation.
We use thin plate regression spline smooths to model the effect of time and its interactions with gender and tone pattern. These splines allow the model to flexibly capture how pitch contours unfold over time across groups and are well suited for continuous, non-linear trajectories such as fundamental frequency (F0) contours. For an introduction to the use of GAMMs in phonetics, see [51]; for mathematical details, see [52].
Thin plate regression splines are also applied to continuous predictors such as duration, word position, and bigram probabilities. Tensor product smooths are used to model interactions between time and continuous predictors, with the ti() directive to isolate the interaction effect from its corresponding main effects. Factor smooths for time are incorporated for tonal context, speaker, and word. These factor smooths (specified as s(time, predictor, bs="fs", m=1)) represent nonlinear random effects of subject and word, and include shrunk estimates of by-subject and by-word intercepts.
@@ 3 RESULTS
The model that we propose for the present data was obtained by fitting a series of increasingly complex GAMM to the dataset. The best-fit model resulting from this exploratory investigation of the Mandarin pitch contours is specified follows:
logF0 $\sim$ genderXtone +